OpenStudy (anonymous):

find an equation of the line containing the given pair of points. (-7,0) & (0,2) y=

OpenStudy (anonymous):

use point-slope formula, y-y1=b(x-x1), and we have: y-(0)=(2/7)(x-(-7). Solving for y: y=(2/7)x

OpenStudy (anonymous):

have a good weekend

OpenStudy (anonymous):

have a good weekend

OpenStudy (anonymous):

thank you and you too.